# TreeCat: Standalone Catalog Engine for Large Data Systems

### v1

@@ Abstract.
With ever increasing volume and heterogeneity of data, advent of new specialized compute engines, and demand for complex use cases, large scale data systems require a performant catalog system that can satisfy diverse needs. We argue that existing solutions, including recent lakehouse storage formats, have fundamental limitations and that there is a strong motivation for a specialized database engine, dedicated to serve as the catalog. We present the design and implementation of TreeCat, a database engine that features a hierarchical data model with a path-based query language, a storage format optimized for efficient range queries and versioning, and a correlated scan operation that enables fast query execution. A key performance challenge is supporting concurrent read and write operations from many different clients while providing strict consistency guarantees. To this end, we present a novel MVOCC (multi-versioned optimistic concurrency control) protocol that guarantees serializable isolation. We conduct a comprehensive experimental evaluation comparing our concurrency control scheme with prior techniques, and evaluating our overall system against Hive Metastore, Delta Lake, and Iceberg.
PVLDB Reference Format:
PVLDB, 19(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 19, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/umddb/treecat.
@@ 1. Introduction
The catalog is an important component of a database management system, responsible for storing and maintaining metadata. It contains various types of information, including the properties of logical entities such as databases, tables, views, triggers, and schemas; mapping of these entities to the underlying physical storage; and access control information, to name a few. Because metadata operations are in the hot path of most database operations, a database management system cannot operate efficiently without a high-performance catalog.
In a monolithic data system, the catalog is implemented as a component tightly coupled with the rest of the system. This design suffices when there are only one or a few data systems under operation. But, in a typical enterprise setting, this is rarely the case. Since no single one-size-fits-all system can meet the requirements of all different use cases, organizations often use many specialized engines, each serving a different function. Because each system maintains its own catalog with limited external access, data cannot be easily shared and ETL (extract-transform-load) pipelines need to be set up to migrate data between different systems. However, with the increasing variety and volume of data, along with the continuing emergence of new systems, maintaining and building these pipelines becomes increasingly cumbersome.
@@
At the same time, there has been a growing trend toward the disaggregation of large-scale data systems into smaller subsystems. This includes the separation of compute and storage layers, the increasing adoption of common table formats (parquet, ORC), and the use of “connectors” to link different systems. These efforts aim to achieve a composable data stack, enabling organizations to store data in a shared storage layer and choose from multiple compute engines. However, as long as each system’s metadata about the “managed” data remains separate, maintaining numerous systems will continue to be challenging. For data management systems to be truly composable, the catalog functionality has to be disaggregated from the rest of the system. This issue is now widely recognized in the data systems industry, and there are multiple active projects, including lakehouse storage formats (Foundation, 2024d, c, a), and, most recently, catalog services (Databricks, 2025, 2024; Sakdeo, 2024; Snowflake, 2024a), mainly driven by the industry. However, as we elaborate in the next section, these systems have architectural designs that impose fundamental limitations on both current and future use cases.
In this context, the catalog itself can be seen as an independent database (sub-)system specialized for handling metadata operations requested by various database engines. As such, the standalone catalog has its own set of functional and systemic requirements, which are not adequately addressed by existing solutions.
@@
In this paper, we first characterize the key requirements for an ideal standalone catalog system. We then present our prototype catalog engine, TreeCat, with the following design choices. (1) Given the hierarchical nature of metadata logical modeling and query access patterns, we adopt a hierarchical data model along with a path navigation query language similar to XPath. (2) For efficient query execution, we employ a storage layout inspired by file systems, implemented using write-optimized key-value stores, and leverage batch correlated scan operations. (3) We use versioned storage to support MVCC (multi-version concurrency control) for improved read performance, as well as version control operations such as time travel queries, clone, and snapshot. (4) We present a novel MVOCC mechanism that combines scan range locking and precision locking techniques. This approach efficiently manages predicate dependencies while minimizing validation costs. Additionally, we adopt a commit-time update technique to avoid aborts caused by conflicts on frequently updated fields, such as statistics. Although we focus on catalogs in this paper, our techniques are more generally applicable in any hierarchical database management system. TreeCat is implemented in about 12000 lines of C++ code, and exposes a gPRC interface to serve remote client requests. We present a comprehensive set of experiments illustrating the significant performance benefits of our new MVOCC protocol. We also compare TreeCat with Delta Lake, HMS, and Iceberg to show that TreeCat supports read operations and commit protocol with significantly lower latency.
@@
2. Background and Prior Work11footnotemark: 1
In traditional RDBMSs and most OLTP systems today, catalog metadata is stored as a predefined set of relations (Hellerstein et al., 2007; PostgreSQL, [n.d.]; MySQL, 2016). Apart from a specialized in-memory cache for low-latency reads, all metadata, including information about namespaces, databases, index, access control, and statistics, use the same underlying storage representation and concurrency control mechanism as the data. Because both metadata operations and data operations are transactional, and the volume of metadata is relatively small, such implementations are sufficient in most cases. However, the tight integration and specialized code make it difficult to extend catalog functionality beyond a single system, or to enrich it to support emerging requirements such as versioning.
With the emergence of data warehouses as primary OLAP systems, different catalog implementations have been developed to handle increasingly decentralized and federated scenarios. While implementations vary across systems, the catalog of a typical data warehouse functions as a specialized subsystem that decouples metadata from data, since the primary data processing infrastructure is not optimized for transactional workloads. Nevertheless, monolithic design persists, as the catalog remains tightly integrated with the rest of the system (Motivala, 2018; Armenatzoglou et al., 2022; Vandiver et al., 2018). For example, Snowflake, a cloud-native data warehouse, stores its metadata in FoundataionDB (Zhou et al., 2021), a distributed transactional key-value store. Many of its features, including automatic data clustering, data pruning during query execution, access control, and zero-copy clone, rely on catalog objects persisted in FoundationDB (Snowflake, 2024b; Motivala, 2018). However, because external systems cannot directly access the metadata, they cannot perform as efficiently or offer the same features when handling tables actively managed by Snowflake.
On the other hand, large-scale data processing engines (Thusoo et al., 2010; Armbrust et al., 2015; Sethi et al., 2019) from the big data community originally emerged with a highly disaggregated architecture, consisting of an independent compute engine, distributed file system, and a metadata service, each designed to be highly scalable. As part of the Hive ecosystem, HMS (Hive Metastore) (Thusoo et al., 2010) was commonly used as the standalone catalog system. Backed by a transactional RDBMS (typically MySQL or PostgreSQL), HMS stores information about databases, tables, access control information, statistics, and partitions (groups of file objects organized by common attribute values). Unfortunately, HMS has a major shortcoming: it does not maintain file-level metadata. Instead, it relies on the underlying storage system (distributed file system or object store) to group files in the same partition under the same directory, essentially using the directory structure as an index. To scan a table, a system must retrieve the paths of relevant partitions from HMS and call list() on each partition directory to obtain the list of files to scan. In environments where data could scale to millions of partitions, the high latency of list() operation of cloud object stores like S3 becomes a serious performance bottleneck (Armbrust et al., 2020). In addition, even retrieving the list of large number of partitions from HMS is slow due to an inefficient Thrift server implementation and suboptimal back-end database schema (Katz, 2021).
As these limitations of HMS became apparent, lakehouse storage formats, namely Delta Lake, Hudi, and Iceberg (Armbrust et al., 2020; Foundation, 2024a, c) emerged as alternatives. While the specific implementations differ, the common idea is to store metadata about file objects as separate metadata files along with the data files in shared storage. By directly fetching and processing these metadata files, the performance bottlenecks of HMS can be avoided. Furthermore, additional file metadata, such as statistics, can be stored in metadata files and utilized for optimizations like file pruning, further improving system performance. Delta Lake uses a data structure called Delta log, which consists of a sequence of delta JSON files (with incrementing numerical IDs) and periodic checkpoint files that store the complete list of file objects at specific snapshots. Iceberg organizes metadata into a hierarchy of files: a metadata file contains a list of manifest list files; a manifest list file contains a list of manifest files; and a manifest file contains a list of file objects. Hudi identifies files by the timestamps at which they are first added and stores a set of indices in HFile format, enabling fast pruning of files based on attributes.
Storing the metadata alongside the data provides additional benefits such as a simplified architecture, good scalability due to the underlying storage system, and speedup from parallel metadata processing. However, this approach has notable drawbacks, which are becoming increasingly problematic as lakehouse systems evolve to support more features. One major drawback is the lack of an advanced concurrency control mechanism, especially multi-table transactions. Another issue is the overhead of operating on file-level metadata, which requires fetching metadata files from shared storage before processing them, resulting in higher latency. Lastly, every lakehouse system still requires a separate catalog system to store the metadata about high-level catalog objects, such as databases and tables. More recently, several efforts have focused on providing these services through Hive-compatible APIs or REST APIs (Databricks, 2025, 2024; Sakdeo, 2024; Snowflake, 2024a). Ironically, contrary to the motivation of replacing HMS with a simpler scheme, lakehouse storage formats are becoming part of even more complex ones, where metadata spans multiple components.
@@ 2. Background and Prior Work11footnotemark: 1
Enterprise data catalogs serve as an organization-level database of metadata about datasets across various systems. There are numerous commercial data catalog software products, including those by Alation, Collibra, Informatica, AWS (Glue), Google (Dataplex), and Microsoft (Azure Data Catalog), as well as open source projects, such as Apache Atlas (Foundation, 2018), Metacat (Netflix, 2024), and DataHub (Authors, 2025a). Unlike operational catalogs, the primary goal of data catalogs is organization-wide data governance, which involves constructing and maintaining metadata stores through data discovery processes (Vahlkamp, 2024; Authors, 2025b; Majumdar and Li, 2018; Services, 2025). However, systems like AWS Glue also offer functions traditionally provided by operational catalogs. Although, the distinctions between them are becoming increasingly blurred, the performance characteristics of the enterprise data catalogs make them unsuitable for use as operational catalogs today.
3. Requirements
Some of the key requirements for a standalone operational catalog are as follows.
Functional Requirements: The key functional requirements for a catalog, as derived from our investigations, are as follows: (1) Manage metadata about numerous logical objects and their relationships, which are fundamentally hierarchical (but not necessarily tree-structured), providing semantic meaning to the underlying data. (2) Support various statistics and clustering information at the level of physical objects, which is increasingly utilized by database engines for cost-based query optimization. (3) Provide a flexible query language, or API, to support different types of queries. (4) Incorporate explicit versioning semantics to support version control operations, such as time travel queries and clone operations. (5) Be extensible, so a large variety of data sets (e.g., non-tabular data) and database engines can be supported.
System Requirements: Supporting the core functionalities described above imposes the following systems-level requirements: (1) Efficiently handle transactional read operations with low latency (fast range queries are especially important). (2) Support frequent insert, delete, and update operations that accompany ETL (extract, transform, load) processes. (3) Implement a scalable concurrency control mechanism with strong ACID guarantees.
Out-of-scope Requirements: Finally, we briefly outline important considerations that are beyond the scope of this paper, but drive our future research: (1) Garbage collection, required to remove obsolete physical files and objects, (2) Availability and fault tolerance, which is essential in a distributed architecture, and (3) User-defined Types and Functions, which are powerful features that further enhance extensibility.
In addition, there are two important issues closely aligned with the catalog, but should be done externally to the catalog. First, maintaining consistency between the data and the metadata is a major emerging challenge. However, it is unclear where that responsibility should lie. While our proposed catalog system does not enforce data-metadata consistency internally, it provides functionality for external systems to manage it. Similarly, access control is a crucial challenge in this space, but needs to be enforced by the layer that manages access to the data itself, which most catalogs (including TreeCat) don’t.
4. High Level Design
@@
In this section, we present the high-level data model and the basic API of TreeCat.
@@ 4.1. Data Model
TreeCat adopts a hierarchical data model. A typical metadata model of a database consists of a nested hierarchy of namespaces, databases, tables, partitions, and file objects. Even in the case of non-tabular data sets, the hierarchical model is suitable because some logical grouping of the underlying data is necessary. The model takes advantage of data locality, where objects within the same logical group are often accessed together. To support various systems, the model should be structurally flexible, of arbitrary depth, and hold objects of different types, as illustrated below:
Figure 2. An example retail database, which shows a typical nested hierarchical structure of a relational database system.
Example 1: Figure 2 illustrates a retail database consisting of multiple tables, which are further partitioned by multiple attribute values. The lowest-level partitions contain physical files and database blocks.
Example 2: Figure 3 illustrates datasets containing some metrics stored in a data lake. One dataset is a data stream of active metric logs ingested in batches. The other dataset is aggregate data stored as a table and is derived from the active log data.
We now proceed with formal definitions of our data model.
@@
      A non-leaf object has a parent and a set of children objects. It is uniquely identified by its path, $/oid_{1}/oid_{2}/../oid_{n}$, where $oid_{n}$ is its object id and the object with the path of $/oid_{1}/oid_{2}/.../oid_{n-1}$ is its parent.
@@
      A leaf object has one or more parents, but does not have any child object. A leaf object with multiple parents has multiple paths. Given one of its paths, $/oid_{1}/oid_{2}/../oid_{n}$, $oid_{n}$ is one of its object ids and the object with the path of $/oid_{1}/oid_{2}/.../oid_{n-1}$ is one of its parents. Unlike non-leaf objects, a leaf object is immutable and can only be added or removed.
@@
    Every change (object addition, removal, or update) is associated with a $vid$ (version id) that determines its globally consistent chronological order.
Some of our design choices deviate from standard practices, so are worth documenting:
  1. (1)
    Because leaf objects often represent physical objects (e.g., Parquet files) that are short-lived and almost never modified, they are immutable, avoiding the storage of past versions of the metadata.
  2. (2)
    Because a leaf object may have multiple non-leaf parents, the hierarchical structure of the data model takes the form of a DAG, which enables sharing datasets without duplicating the underlying physical objects or their metadata.
  3. (3)
    TreeCat provides versioning semantics due to its high utility. But, considering practical functionality, we only provide snapshot operation and clone operation, as described below.
Figure 3. An example metrics data sets, consisting of a data stream and a table.
4.2. Basic API and Query Language
Next, we list a basic API that TreeCat supports.
  • startTransaction(mode): Starts a new transaction in the specified mode. If the mode is read-only, it returns the latest $vid$ at which a snapshot of the catalog can be read. If the mode is read-write, it creates a new transaction session and returns $transaction\_id$ and $read\_vid$.
  • executeQuery(query, optional transaction_id, optional vid)): Executes a read query on the catalog using the query language, defined below.
  • commit(transaction_id, write_set): Validates and commits the transaction of identified by $transaction\_id$, applying modifications specified by the given $write\_set$.
  • snapshot(snapshot_name, vid): Creates an immutable snapshot of the catalog data at $vid$ and names it $snapshot\_name$.
  • clone(src_path, dest_path, optional vid): Clones/copies the object at $src\_path$ and all of its descendants to the $dest\_path$. Note that this is a metadata-level operation, and does not involve cloning the physical objects.

### v2

@@ Abstract.
With ever-increasing volume and heterogeneity of data, advent of new specialized compute engines, and demand for complex use cases, large-scale data systems require a performant catalog system that can satisfy diverse needs. We argue that existing solutions, including recent lakehouse storage formats, have fundamental limitations and that there is a strong motivation for a specialized database engine, dedicated to serve as the catalog. We present the design and implementation of TreeCat, a database engine that features a hierarchical data model with a path-based query language, a storage format optimized for efficient range queries and versioning, and a correlated scan operation that enables fast query execution. A key performance challenge is supporting concurrent read and write operations from many different clients while providing strict consistency guarantees. To this end, we present a novel MVOCC (multi-versioned optimistic concurrency control) protocol that guarantees serializable isolation. We conduct a comprehensive experimental evaluation comparing our concurrency control scheme with prior techniques, and evaluating our overall system against Hive Metastore, Delta Lake, and Iceberg.
@@ 1. Introduction
The catalog is an important component of a database management system that is responsible for storing and maintaining metadata such as the properties of logical entities like databases, tables, views, triggers, and schemas; mapping of these entities to the underlying physical storage; and access control information. Metadata operations are in the hot path of most, if not all, database operations. Take, for example, the query execution process of a relational database. The table schema is necessary to first semantically validate the query. Once the query is validated, the query planner uses various metadata, including table schema, storage information, and statistics, to optimize and build a physical query plan that is executed by the execution engine (which usually needs access to the metadata as well). Without a high-performance catalog that provides fast access to metadata, a database management system cannot operate efficiently.
@@
In a monolithic data system, the catalog is typically tightly coupled with the rest of the system, which works for one or a few data systems. But in a typical enterprise setting, organizations often use many specialized engines, each serving a different function, since no single one-size-fits-all system can meet the requirements of all use cases. Meanwhile, there has been a growing trend toward disaggregation of large-scale data systems into smaller subsystems. This includes the separation of compute and storage layers, the increasing adoption of common table formats (parquet, ORC), and the use of “connectors” to link different systems. These efforts aim to achieve a composable data stack, enabling organizations to store data in a shared storage layer and choose from multiple compute engines. However, as long as each system maintains its own catalog such that the metadata about the “managed” data remain separate, different systems cannot easily share data even if it is stored in the same storage system. For data management systems to be truly composable, the catalog functionality has to be disaggregated from the rest of the system. This issue is now widely recognized in the data systems industry, and there are multiple active projects, including lakehouse storage formats (Foundation, 2025f, c, a), and, most recently, catalog services (Databricks, 2025a, b; Sakdeo, 2024; Snowflake, 2024a), mainly driven by the industry.
We further argue that not only should the catalog functionality exist as a standalone service, but also that it should be unified and not span multiple (semi-)autonomous systems or storage modules. As we show, unified architecture not only leads to significantly reduced latencies, but also reduces duplication of metadata across different data engines, and enables multi-table transactions as well as other complex operations on metadata without sacrificing consistency guarantees. A unified design also makes it possible to build a specialized engine, as we develop in this work, that can meet the specific requirements of a catalog in terms of both functionality and performance.
@@
To this end, the proposed catalog itself can be seen as an independent database (sub-)system specialized for handling metadata operations. We characterize the primary use cases and present our prototype catalog engine, TreeCat, with the following design choices. (1) Given the hierarchical nature of metadata logical modeling and query access patterns, we adopt a hierarchical data model along with a path navigation query language. (2) For efficient query execution, we employ a storage layout inspired by file systems, implemented using write-optimized key-value stores, and leverage batch correlated scan operations. (3) We use versioned storage to support MVCC (multi-version concurrency control) for improved read performance, as well as version control operations such as time travel queries, clone, and snapshot. (4) We present a novel MVOCC mechanism that combines scan range locking and precision locking techniques. This approach efficiently manages predicate dependencies while minimizing validation costs. Additionally, we adopt a commit-time update technique to avoid aborts caused by conflicts on frequently updated fields, such as statistics. Although we focus on catalogs in this paper, some of our techniques are more generally applicable in any hierarchical database management system. TreeCat is implemented in about 12000 lines of C++ code, and exposes a gPRC interface to serve remote client requests. We present a comprehensive set of experiments illustrating the significant performance benefits of our new MVOCC protocol. We also show that TreeCat outperforms other state-of-the-art systems, including Delta Lake, HMS, and Iceberg.
@@
2. Background and Prior Work
In this section, we categorize and classify existing catalog systems and highlight their shortcomings, which motivate the key requirements and features of TreeCat. The discussion of prior work related to more specific design and implementation details is interspersed in other relevant parts of the paper for better context.
@@ 2. Background and Prior Work
We first distinguish between data catalogs and operational catalogs. Data catalogs serve as an organization-level database of metadata about data assets across various systems. They can be viewed as an “information retrieval system” for data assets. The main goal is to make the data assets of an organization more manageable and accessible, providing functions such as keyword search and retrieval of important contextual information. Data catalogs are typically implemented as federated systems that are updated asynchronously through data discovery processes or metadata ingestion pipelines (Vahlkamp, 2024; DataHub, 2025b; Majumdar and Li, 2018; Services, 2025). Because they do not directly serve database engines, they generally lack strict consistency or performance guarantees. There are numerous commercial data catalog software products, including those by Alation, Collibra, Informatica, AWS (Glue), Google (Dataplex), and Microsoft (Azure Data Catalog), as well as open source projects, such as Apache Atlas (Foundation, 2019), Metacat (Netflix, 2025), and DataHub (DataHub, 2025a). Operational catalogs, on the other hand, store metadata that are directly accessed by database engines during updates, query planning, and query execution. Although operational catalogs can be a part of, or serve as the main data source for, data catalogs, their primary functions are clearly different. In this paper, we focus exclusively on operational catalogs.
Catalogs can be either embedded or standalone. An embedded catalog is integrated within a single data system, and external access to its metadata is restricted to the interfaces provided by the parent system. A standalone catalog, on the other hand, exists as an independent service that is accessible by any compatible data system via a standard interface. Catalogs can be further differentiated based on their architectural designs. A catalog with a modular architecture is composed of components spread across multiple autonomous services and systems. On the other hand, a catalog with a unified architecture consists of tightly integrated components forming a single system. It is important to note that a unified system is not necessarily centralized; it can have a distributed architecture for scalability.
Figure 2. Classification of different catalog systems.
Traditional RDBMSs, including both transactional and analytical systems, have embedded catalogs with unified architectures. In transactional RDBMSs, catalog metadata is stored as a predefined set of relations (Hellerstein et al., 2007; PostgreSQL, 2025; MySQL, 2016). Apart from a specialized in-memory cache for low-latency reads, all metadata use the same underlying storage representation and concurrency control mechanism as the data. Analytical RDBMSs and data warehouses have different implementations designed to handle increasingly decentralized scenarios. The catalog of a typical data warehouse functions as a specialized subsystem that decouples metadata from data, since the primary data processing infrastructure is not optimized for transactional workloads. For example, Snowflake, a cloud-native data warehouse, stores its metadata in FoundationDB (Zhou et al., 2021), a distributed transactional key-value store. Many of its features, including automatic data clustering, data pruning during query execution, access control, and zero-copy clone, rely on catalog objects persisted in FoundationDB (Snowflake, 2024b; Motivala, 2018). Regardless of system performance, tight integration with the rest of the system makes it difficult to extend the catalog functionality beyond a single system (Motivala, 2018; Armenatzoglou et al., 2022; Vandiver et al., 2018).
Large-scale data processing engines (Thusoo et al., 2010; Armbrust et al., 2015; Sethi et al., 2019) from the big data community use standalone catalogs with modular architectures. These systems originally emerged with a highly disaggregated architecture consisting of an independent compute engine, a distributed storage system, and a metadata service, each designed to be highly scalable. As part of the Hive ecosystem, HMS (Hive Metastore) (Thusoo et al., 2010) was commonly used as the standalone catalog system. Backed by a transactional RDBMS (typically MySQL or PostgreSQL), HMS stores information about databases, tables, access control information, statistics, and partitions (groups of file objects organized by common attribute values). However, HMS does not store file-level metadata in the back-end RDBMS. Instead, it uses the shared storage system (distributed file system or object store) as an index on the files by grouping them in the same partition under the same directory (hence, it is modular). To scan a table, HMS must retrieve the paths of relevant partitions from the back-end RDBMS and call list() on each partition directory to obtain the list of files to scan. In an environment where a single dataset contains over a million files, the high latency of list() operation of cloud object stores like S3 becomes a serious performance bottleneck (Armbrust et al., 2020). In addition, even retrieving the list of a large number of partitions from HMS is slow due to an inefficient Thrift server implementation and suboptimal back-end database schema (Katz, 2021).
As the limitations of HMS became clear, lakehouse storage formats, namely Delta Lake, Hudi, and Iceberg (Armbrust et al., 2020; Foundation, 2025a, c) emerged as alternatives. The specific implementations differ, but the common idea is to store metadata about file objects as separate metadata files along with the data files in shared storage. By directly fetching and processing these metadata files, the performance bottlenecks of HMS can be avoided. Furthermore, additional file metadata, such as statistics, can be stored in metadata files and utilized for optimizations like file pruning, further improving system performance. Delta Lake uses a data structure called Delta log, which consists of a sequence of delta JSON files (with incrementing numerical IDs) and periodic checkpoint files (Armbrust et al., 2020). Iceberg organizes metadata into a hierarchy of files: a metadata file contains a list of manifest list files; a manifest list file contains a list of manifest files; and a manifest file contains a list of file objects (Foundation, 2025e). Hudi identifies files by the timestamps at which they are first added and stores metadata in HFile format, enabling fast file pruning and listing (Foundation, 2022).
Storing the metadata alongside the data has benefits such as a simplified architecture, good scalability due to the shared storage system, and speedup from parallel metadata processing. However, the notion that the adoption of the lakehouse storage format results in a simpler architecture than HMS is only partially true. A separate system must still be deployed to manage the metadata about high-level catalog objects, such as databases and tables; lakehouse storage formats simply offload most of the catalog functions to the shared storage system. More recently, several efforts have focused on providing these services through Hive-compatible APIs or REST APIs (Databricks, 2025a, b; Sakdeo, 2024; Snowflake, 2024a). Due to the modular architecture, it is still quite difficult to maintain consistency of the metadata. As discussed in Section 5.4, no lakehouse storage format currently supports multi-table transactions. Lastly, every metadata operation involves reading and writing metadata files in the shared storage system, which has non-trivial overhead. The high latency of these operations can limit the write throughput under high contention (Jain et al., 2023).
3. High-Level Design
@@
In this section, we present the high-level data model and the basic API of TreeCat. We also characterize the primary use cases with concrete examples.
@@ 3.1. Data Model
A catalog needs to manage metadata about numerous logical objects and their relationships, which are almost always organized into nested logical groups, naturally forming a multi-level hierarchy. It also needs to maintain arbitrary statistics and clustering information (for cost-based optimizations) and must incorporate explicit versioning semantics to support operations such as time-travel and cloning. Further, it should support a sufficiently powerful query language and/or API so that the clients don’t end up replicating the functionality, while being flexible and extensible enough to support diverse data sets and database engines. This naturally leads us to adopt a hierarchical data model for TreeCat. We use an example Retail database shown in Figure 3 to illustrate this point.
Figure 3. An example retail database, which contains heterogeneous assets, including relational data and ML models.
Example 1: The retail database primarily contains relational data, but also includes other types of assets, such as ML models. Furthermore, different types of assets have different nested substructures. While an ML model consists of different versions, each identified by a tag, data in a relational table are partitioned by a set of attribute values. Leaf objects correspond to physical data objects in the storage system.
We further note that: (1) Each group has common contextual information that is important for handling the constituent entities. (2) A predicate is often evaluated against such information, allowing entire groups of irrelevant objects to be pruned out during the data retrieval process. (3) There is data locality where entities in the same group are likely to be accessed together. These points are illustrated by the example use case that we present later. We now proceed with formal definitions of TreeCat’s data model.
@@
      A non-leaf object has a parent and a set of children objects. It is uniquely identified by its path, $oid_{1}/oid_{2}/../oid_{n}$, where $oid_{n}$ is its object id and the object with the path of $oid_{1}/oid_{2}/.../oid_{n-1}$ is its parent.
@@
      A leaf object has one or more parents, but does not have any child object. A leaf object with multiple parents has multiple paths. Given one of its paths, $oid_{1}/oid_{2}/../oid_{n}$, $oid_{n}$ is one of its object ids and the object with the path of $oid_{1}/oid_{2}/.../oid_{n-1}$ is one of its parents. Unlike non-leaf objects, a leaf object is immutable and can only be added or removed.
@@
    Every change (object addition, removal, or update) is associated with a vidvid (version id) that determines its globally consistent chronological order.
We highlight a few important design decisions that deviate from standard practices. (1) Immutable leaf objects: Leaf objects, which usually correspond to physical data objects (e.g., Parquet files), are treated as immutable to enable seamless versioning and sharing — updates can be handled as delete‐plus‐insert if needed. (2) Shared parents: Allowing leaf objects to have multiple parents lets datasets be shared, cloned or branched, without duplicating physical files or metadata. (3) Global chronological versions: A single, consistent timeline underlies our concurrency control and powers version‐control features such as time‐travel queries, snapshots, and clones.
Note that, while Figure 3 shows one possible instantiation of metadata schema, the engine imposes no fixed metadata model—users choose whatever schema best fits their use case.
3.2. Query Language